OpenStudy (anonymous):

find all real numbers if any that are fixed points for the givem functions. F(t)=t^2-t-1

OpenStudy (anonymous):

do you know what a fixed point means?

OpenStudy (anonymous):

can you show your work

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

n

OpenStudy (anonymous):

alright, tell me.

OpenStudy (anonymous):

It is the point that intersects

OpenStudy (anonymous):

you have to be clearer than that,

OpenStudy (anonymous):

a point on the x or y axis

OpenStudy (anonymous):

or when two lines intersect

OpenStudy (anonymous):

okay. a fixed point is a point such that F(n) = n.

OpenStudy (anonymous):

that is were I start to get confused

OpenStudy (anonymous):

So F(t)=t

OpenStudy (anonymous):

yes. we know that F(t) = t^2-t-1. therefore t^2-t-1 = t. solve for t and you have your points.

OpenStudy (anonymous):

can you walk me through that?

OpenStudy (anonymous):

(t-t)(t-1)

OpenStudy (anonymous):

it is an equation. what is your right hand side?

OpenStudy (anonymous):

t^2-t-1

OpenStudy (anonymous):

no. you had t^2-t-1 = t that is an equation. subtract t from both sides. what do you get?

OpenStudy (anonymous):

0=t^2-2t-1

OpenStudy (anonymous):

0=(t+1)^2